OpenStudy (anonymous):

The temperature of a chemical reaction ranges between 40°C and 180°C. The temperature is at its lowest point when t = 0 and completes one cycle over a 12-hour period. What is a sine function that would model this reaction?

OpenStudy (anonymous):

Please help! I'll fan and medal

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

\[f(t)=70\sin((\pi/6)t - \pi/2) +110\]